OpenStudy (idealist10):

Find the general solution of y'+(1/x-1)y=-2/x.

OpenStudy (cydney_morgan):

@iambatman Knows how to do this, I don't Sorry :C

OpenStudy (idealist10):

@iambatman

OpenStudy (anonymous):

Differential equations? What have you done so far?

OpenStudy (idealist10):

Yes, I tried to do integration by parts but it only complicated the problem more.

OpenStudy (idealist10):

Please help me. I've stucked on this problem for days.

OpenStudy (anonymous):

\[y'+\frac{ 1 }{ x-1 }y=-\frac{ 2 }{ x }\] right?

OpenStudy (idealist10):

No, the middle part is ((1/x)-1)y

OpenStudy (anonymous):

\[y'+(\frac{ 1 }{ x }-1)y=-\frac{ 2 }{ x }\]

OpenStudy (idealist10):

Yes, please continue.

OpenStudy (cydney_morgan):

He is not going to give you the answer, he is going to help you and teach you how to get the answer yourself ♩ヽ(・ω・ヽ*)♬

OpenStudy (idealist10):

Okay.

OpenStudy (anonymous):

\[\frac{ dy }{ dx }+\frac{ (-x+1)y }{ x }=-\frac{ 2 }{ x }\] solving as a linear equation right?

OpenStudy (idealist10):

Yes.

OpenStudy (anonymous):

I was about to call ganeshie as back up but he's already here :3

ganeshie8 (ganeshie8):

hey are you the same Idealist that kept me waiting the other day :o

OpenStudy (idealist10):

Yes, the other website doesn't work so I've created a new account.

OpenStudy (anonymous):

\[\huge u = e ^{\int\limits \frac{ -x+1 }{ x }dx}=e ^{-x}x?\] so \[\frac{x* \frac{ dy }{ dx } }{ e^x }+\frac{ (-x+1)y }{ e^x }=-\frac{ 2 }{ e^x }\] multiplying both sides by u, Idk if this is right though, ganeshie knows more...

ganeshie8 (ganeshie8):

Okay haha! i think iambatman is already helping you, ima just watch how you work this... :) this is really an interesting problem if you're doing linear DEs for the first time !

OpenStudy (anonymous):

Nuuu go ahead ganeshie :3

OpenStudy (cydney_morgan):

So many smart people (づ。◕‿‿◕。)づ

ganeshie8 (ganeshie8):

that looks perfect to me !

ganeshie8 (ganeshie8):

the whole idea of multiplying \(\large u\) both sides is to write the left side as \(\large (uy)'\) , @Idealist10 do you see how to apply the product rule in reverse on left side ?

OpenStudy (idealist10):

So the integrating factor is (e^-x)(x)?

OpenStudy (anonymous):

Substitute \[e ^{-x}(1-x) = \frac{ d }{ dx }\left( \frac{ x }{ e^x } \right)\]

OpenStudy (idealist10):

In the product rule (uy)', since we already found u, the integrating factor, what's y?

OpenStudy (anonymous):

Mhm, Well you have \[\frac{ x*\frac{ dy }{ dx } }{ e^x }+\frac{ d }{ dx }\left( \frac{ x }{ e^x } \right)y=-\frac{ 2 }{ e^x }\] So now you want to apply reverse product rule, \[\frac{ d }{ dx }\left( \frac{ xy }{ e^x } \right)=-\frac{ 2 }{ e^x }\] Now you can integrate... \[\int\limits \frac{ d }{ dx }\left( \frac{ xy }{ e^x } \right)dx=\int\limits \left( -\frac{ 2 }{ e^x } \right)dx\] the reverse product rule is \[g \frac{ df }{ dx }+f \frac{ dg }{ dx }=\frac{ d }{ dx }(fg)\] Once again thought I'm not entirely sure if this is 100% correct haha.

ganeshie8 (ganeshie8):

thats looks good to me, @Idealist10 : after some practice, you can blindly write the left side as \(\large (\color{Red}{u}y)'\)... to your earlier question, \(y\) is just \(y\), left side becomes \(\large (\color{red}{\dfrac{x}{e^x}}y)'\) (shown as dy/dx notation in above reply)

OpenStudy (anonymous):

Yeah sorry lost connection

OpenStudy (anonymous):

Does this make sense?

OpenStudy (idealist10):

Thank you so much for the help. I'm sorry it took long to understand. I just find this topic hard.

OpenStudy (anonymous):

No worries, no reason to apologize for that :). I actually like it when people take their time and try to learn subjects ^.^